OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3?

OpenStudy (anonymous):

@Fanduekisses <3<3

OpenStudy (fanduekisses):

You need to use the point slope formula

OpenStudy (fanduekisses):

\[y-y_{1}=m(x-x_{1})\]

OpenStudy (fanduekisses):

Plug in what you were given

OpenStudy (anonymous):

Jesus thats confusing oh dear lord

OpenStudy (anonymous):

m = rise/run ??

OpenStudy (fanduekisses):

yes, m = slope which in this case is -3

OpenStudy (anonymous):

What about 4, -2?? I'm sorry this is very confusing >.<

OpenStudy (fanduekisses):

remember (x,y) so x=4 and y = -2

OpenStudy (fanduekisses):

but plug them in the\[x_{1} and y _{1} \]

OpenStudy (fanduekisses):

\[(y--2)=-3(x-4) \]

OpenStudy (fanduekisses):

Simply/ distribute

OpenStudy (fanduekisses):

then get it in standard form which is in Ax + By + C = 0

OpenStudy (fanduekisses):

After simplifying you should have: \[y+2=-3x+4\]

OpenStudy (fanduekisses):

then get it in standard form

OpenStudy (anonymous):

What would be standard form exactly ?

OpenStudy (anonymous):

???

OpenStudy (fanduekisses):

No, the standard form is Ax+Bx= C